# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v1

@@ 3.1. Relative stability conditions
Recall that each ordinary GM fourfold $X$ has a canonical quadric, which is smooth when $X$ is general (cf. [59, Section 2.1, Remark 2.2]). More generally, for a family of ordinary GM fourfolds, there is a family of canonical quadrics (cf. [59, Section 5.1]).
@@
This follows from [59, Proposition 5.6]. ∎
@@ 3.3. Lagrangian subvarieties in moduli spaces
For a GM threefold or fivefold $Y$, there is an involution $T_{Y}$ on $\mathcal{K}u(Y)$ (cf. [24, Lemma 4.3]), which acts trivially on $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(Y))$. Therefore, by Theorem 2.8, $T_{Y}$ induces an involution on the moduli space
@@
For a GM fourfold or sixfold $X$, there is also an involution $T_{X}$ on $\mathcal{K}u(X)$ as in [24, Lemma 4.3], which acts trivially on the sublattice $A_{1}^{\oplus 2}\subset\widetilde{\mathrm{H}}(\mathcal{K}u(X),\mathbb{Z})$. The following result is proved in [60]:
Theorem 3.6 ([60, Theorem 1.6]).
@@
Therefore, one can construct a Lagrangian submanifold $\mathrm{Fix}(\tau_{X})$ for each moduli space $M_{\sigma_{X}}^{X}(a,b)$. However, as shown in [24, Corollary 6.8], $\mathrm{Fix}(\tau_{X})$ is a Lagrangian constant cycle subvariety, hence it is expected to be rigid.
@@ 3.4. Bayer–Macrì divisors
for any morphism $C\to\mathcal{M}^{Y}_{\sigma_{Y}}(a,b)$ from an integral curve $C$, where $\mathcal{E}_{C}$ is the associated family and $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}$ is the base change of $T_{Y}$ [^5] The functor $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}\colon\mathrm{D}^{\mathrm{b}}(Y_{C})\to\mathrm{D}^{\mathrm{b}}(Y_{C})$ is defined by the pull-back of the kernel of the composition $\mathrm{D}^{\mathrm{b}}(Y)\xrightarrow{\mathrm{pr}_{Y}}\mathcal{K}u(Y)\hookrightarrow\mathrm{D}^{\mathrm{b}}(Y)$ to $\mathrm{D}^{\mathrm{b}}(Y_{C}\times_{C}Y_{C})$.. Note that $\mathcal{E}_{C}\in\mathrm{D}^{\mathrm{b}}(X_{C})$ by [8, Lemma 8.3(1)]. From [40, Lemma 2.41], we obtain an isomorphism $p_{Y*}(T_{Y_{C}}(\mathcal{E}_{C}))\cong T_{Y}(p_{Y*}(\mathcal{E}_{C}))$. This implies
@@
as $T_{Y}$ acts trivially on $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(Y))$ by [24, Lemma 4.10].
@@
then (11) follows from the base change of [24, Equation (35)]. This completes the proof of (6). Finally, the ampleness of $\ell_{\sigma_{Y}}$ follows from the facts that $\ell_{\sigma_{X}}$ is ample (cf. [62, Theorem 3.1]) and the morphism $r$ is finite according to Theorem 3.7. ∎
@@
Similarly, using [24, Equation (40)] instead of [24, Equation (35)], we have the following result for GM fivefolds.
@@ 4.1.2. Atomic objects
Such a vector $\tilde{v}(E)\in\tilde{\mathrm{H}}(X,\mathbb{Q})$ is called an extended Mukai vector of $E$ (cf. [6, Definition 4.16] and [5, Proposition 3.3]). In [49, Section 6], the $1$-dimensional subspace of $\tilde{\mathrm{H}}(X,\mathbb{Q})$ spanned by $\tilde{v}(E)$ is called the LLV line of $E$.
@@
By our assumption, $\mathcal{X}_{b}$ is smooth, hence $\mathcal{X}_{b}\hookrightarrow\mathcal{X}$ is a locally complete intersection. Moreover, $\mathrm{codim}_{\mathcal{Y}}(\mathcal{Y}_{b})=\mathrm{codim}_{\mathcal{X}}(\mathcal{X}_{b})=\dim B$. Then using [40, Corollary 2.27], we obtain
@@
As in [5] and [49, Lemma 6.10], for a sheaf $E\in\operatorname{\mathrm{Coh}}(X)$, we have
@@ 4.2. Atomic sheaves supported on Lagrangian subvarieties
By [49, Theorem 6.24] and [5, Proposition 3.11], each torsion atomic sheaf is supported on points or Lagrangian subvarieties. We make the following definition.
@@
is the projection. Since $\chi_{i_{*}\operatorname{\mathcal{O}}_{L}}$ is of rank one, it follows that $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ has rank at most one for each $t=1,2$. Then the result follows from the non-triviality of $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ as in the arguments in [49, Lemma 3.9, Remark 3.10] and [5, Section 7.2]. ∎
@@ 4.3. A criterion
To this end, we only need to generalize [5, Proposition 7.2] to our setting. Indeed. the argument there applies analogously without changes. We denote $\mathrm{ker}(\chi^{\mathrm{coh}}_{i_{*}[L]})$ by $K$. Because the image of $i$ is Lagrangian, we see $\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X})\subset K$. Moreover, by the existence of the $\mathfrak{sl}_{2}$-triple $(e_{\omega},h_{\omega},\Lambda_{\omega})\subset\mathfrak{g}(X)_{\mathbb{C}}$ (cf. [65, Section 2]) where $\omega\in\mathrm{H}^{0}(X,\Omega_{X}^{2})$ is the holomorphic symplectic 2-form, the same calculation as in [5, Proposition 7.2] gives $\mathrm{H}^{0}(X,\wedge^{2}\mathcal{T}_{X})\subset K$.
@@
$i_{*}[L]\cup-\colon\mathrm{H}^{2}(X,\mathbb{Q})\xrightarrow{i^{*}}\mathrm{H}^{2}(X,\mathbb{Q})\xrightarrow{i_{*}}\mathrm{H}^{2n+2}(X,\mathbb{Q}).$
@@
The proof of [69, Lemma 1.5] adapts to our situation without changes. It suffices to prove this for cohomology with real coefficients. By the adjunction formula, we have an inclusion $\ker(i^{*})\subset\ker(i_{*}[L]\cup-)$. To show the reverse inclusion, as in the proof of [69, Lemma 1.5], we choose a Kähler class $\lambda\in\mathrm{H}^{2}(X,\mathbb{R})$ such that $i^{*}\lambda$ is also a Kähler class. Now it suffices to prove that the intersection form
@@
is non-degenerate after restricting it to $\mathrm{im}(i^{*})\subset\mathrm{H}^{1,1}_{\mathbb{R}}(L)$. Then the result follows from the Hodge index theorem as in [69, Lemma 1.5] since $i^{*}\lambda$ is a Kähler class. ∎
@@
Let $\omega$ be the holomorphic symplectic $2$-form of $X$. We denote by $\mathcal{N}^{\vee}_{L/X}$ the $-1$ degree cohomology sheaf of the cotangent complex $\mathbb{L}_{L/X}$ of $i$. When $i$ is a closed embedding, $\mathcal{N}^{\vee}_{L/X}$ is exactly the conormal bundle of $L$ in $X$. Since $i$ is a local complete intersection morphism by [64, Tag 0E9K], $\mathcal{N}^{\vee}_{L/X}$ is a bundle (cf. [64, Tag 0FK2] and [64, Tag 068B]). Using the standard exact triangle of cotangent complexes by [64, Tag 0E44], we obtain exact sequences
@@
As $\omega$ is a holomorphic symplectic $2$-form, it induces an isomorphism $\omega\colon\mathcal{T}_{X}\xrightarrow{\cong}\Omega_{X}$. To adapt the argument in [5, Lemma 7.4], it suffices to prove that there exists an isomorphism $\mathcal{N}_{L/X}\cong\Omega_{L}$. According to [31, Remark 4.4], there is a natural map
@@ 5. Atomic sheaves on Bridgeland moduli spaces
In this section, we aim to construct atomic sheaves on $M^{X}_{\sigma_{X}}(a,b)$ for a GM fourfold or sixfold $X$ via two different methods. The main results in this section are Theorem 5.5 and Theorem 5.8.
@@ 5.2. Fixed loci of involutions are $1$-obstructed
First, we assume that $X$ is non-Hodge-special. As in [60, Section 8.5], up to an equivalence of Kuznetsov components, it is enough to prove the statement when $X$ is a non-Hodge-special special GM fourfold. Let $j\colon Y\hookrightarrow X$ be the branch divisor of $X$, which is an ordinary GM threefold. By [24, Proposition 5.3], for any Serre-invariant stability condition $\sigma_{Y}$ on $\mathcal{K}u(Y)$, the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism
@@
Next, we assume that $X$ is an arbitrary smooth GM fourfold. By [59, Theorem 4.18], we may assume that $X$ is general in the sense that the canonical quadric surface in $X$ is smooth (cf. [59, Remark 2.2]). Therefore, using Theorem 3.2, we can find a smooth connected quasi-projective curve $C$ and a family $\mathcal{X}\to C$ of GM fourfolds with a full triangulated subcategory $\mathcal{K}u(\mathcal{X})\subset\mathrm{D}^{\mathrm{b}}(\mathcal{X})$ such that $\mathcal{K}u(\mathcal{X})_{c}\simeq\mathcal{K}u(\mathcal{X}_{c})$ for each $c\in C$. Additionally,
@@
From [12, Lemma 5.9], there is an auto-equivalence $T_{\mathcal{X}}$ on $\mathcal{K}u(\mathcal{X})$ such that $(T_{\mathcal{X}})|_{c}\simeq T_{\mathcal{X}_{c}}$ for each $c\in C$. By [12, Proposition 5.7] and the proof of [60, Theorem 8.2], such equivalence induces an automorphism $\tau_{\mathcal{X}}\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$, which is an involution since it is an involution at each fiber over $C$ and $M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ is smooth. Therefore, the fixed locus $\mathrm{Fix}(\tau_{\mathcal{X}})$ is smooth projective over $C$ and satisfies $(\mathrm{Fix}(\tau_{\mathcal{X}}))|_{c}\cong\mathrm{Fix}(\tau_{\mathcal{X}_{c}})$ for each $c$. By Proposition 5.1, the restriction map
@@
If $Y$ is an ordinary GM fivefold, by [39, Corollary 6.5], we can find an ordinary GM threefold $Y^{\prime}$ with an equivalence $\mathcal{K}u(Y)\simeq\mathcal{K}u(Y^{\prime})$. Thus, we only need to prove the statement when $Y$ is an ordinary GM threefold. As $Y$ is ordinary, it is the branch divisor of a special GM fourfold $X$.
Let $\sigma_{X}\in\operatorname{Stab}^{\circ}(\mathcal{K}u(X))$ be a generic stability condition on $\mathcal{K}u(X)$ with $Z_{\sigma_{X}}(\Lambda_{i})=Z_{\sigma_{Y}}(\lambda_{i})$ for each $i=1,2$. By [24, Proposition 5.3], the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism

### v2

@@ 3.1. Relative stability conditions
Recall that each ordinary GM fourfold $X$ has a canonical quadric, which is smooth when $X$ is general (cf. [66, Section 2.1, Remark 2.2]). More generally, for a family of ordinary GM fourfolds, there is a family of canonical quadrics (cf. [66, Section 5.1]).
@@
This follows from [66, Proposition 5.6]. ∎
@@ 3.3. Lagrangian subvarieties in moduli spaces
For a GM threefold or fivefold $Y$, there is an involution $T_{Y}$ on $\mathcal{K}u(Y)$ (cf. [25, Lemma 4.3]), which acts trivially on $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(Y))$. Therefore, by Theorem 2.8, $T_{Y}$ induces an involution on the moduli space
@@
For a GM fourfold or sixfold $X$, there is also an involution $T_{X}$ on $\mathcal{K}u(X)$ as in [25, Lemma 4.3], which acts trivially on the sublattice $A_{1}^{\oplus 2}\subset\widetilde{\mathrm{H}}(\mathcal{K}u(X),\mathbb{Z})$. The following result is proved in [67]:
Theorem 3.6 ([67, Theorem 1.6]).
@@
Therefore, one can construct a Lagrangian submanifold $\mathrm{Fix}(\tau_{X})$ for each moduli space $M_{\sigma_{X}}^{X}(a,b)$. However, as shown in [25, Corollary 6.8], $\mathrm{Fix}(\tau_{X})$ is a Lagrangian constant cycle subvariety, hence it is expected to be rigid.
@@
The following lemma is an easy consequence of the results in [47]. The proof is similar to that of [47, Theorem 8.2].
Lemma 3.8.
Let $X$ be a general GM fourfold and $j\colon Y\hookrightarrow X$ be a general hyperplane section. Then, for any pair of coprime integers $a,b$ and Serre-invariant stability condition $\sigma_{Y}$ on $\mathcal{K}u(Y)$, there exists a $\sigma_{Y}$-stable object $E\in\mathcal{K}u(Y)$ of class $a\lambda_{1}+b\lambda_{2}$ such that $T_{X}(\mathrm{pr}_{X}(j_{*}E))\neq\mathrm{pr}_{X}(j_{*}E)$.
Proof.
We do induction on $a^{2}+b^{2}$. If $a^{2}+b^{2}\leq 2$, the statement follows from [37] and [25, Appendix B]. Assume that the statement holds whenever $a^{2}+b^{2}<n$ and $n\geq 3$. Now, we will prove the statement for $v:=a\lambda_{1}+b\lambda_{2}$ with $a^{2}+b^{2}=n$.
Let $v_{-}$ and $v_{+}$ be the decomposition of $v$ as in [47, Proposition 2.13]. By the induction hypothesis, we can find $\sigma_{Y}$-stable objects $E_{-},E_{+}\in\mathcal{A}_{\sigma_{Y}}$ such that $T_{X}(\mathrm{pr}_{X}(j_{*}E_{\pm}))\neq\mathrm{pr}_{X}(j_{*}E_{\pm})$. We can assume that $\phi_{\sigma_{Y}}(E_{-})<\phi_{\sigma_{Y}}(E_{+})$. Then, using [47, Lemma 2.12], we can find a $\sigma_{Y}$-stable object $E\in\mathcal{A}_{\sigma_{Y}}$ with $[E]=v$ such that there is an exact sequence $0\to E_{-}\to E\to E_{+}\to 0$ in $\mathcal{A}_{\sigma_{Y}}$.
We claim that $T_{X}(\mathrm{pr}_{X}(j_{*}E))\neq\mathrm{pr}_{X}(j_{*}E)$. Indeed, as $Y$ is general, we have $T_{Y}(E)\neq E$ and $T_{Y}(E_{\pm})\neq E_{\pm}$. If $T_{X}(\mathrm{pr}_{X}(j_{*}E))=\mathrm{pr}_{X}(j_{*}E)$, then by [25, Lemma 5.2] we get $j^{*}\mathrm{pr}_{X}(j_{*}E)\cong E\oplus T_{Y}(E)$. On the other hand, we have the following exact sequence in $\mathcal{A}_{\sigma_{Y}}$
$0\to j^{*}\mathrm{pr}_{X}(j_{*}E_{-})\to j^{*}\mathrm{pr}_{X}(j_{*}E)\to j^{*}\mathrm{pr}_{X}(j_{*}E_{+})\to 0$
This implies that the exact sequence in [25, Lemma 4.10(1)] associated with $j^{*}\mathrm{pr}_{X}(j_{*}E_{\pm})$ splits, which contradicts $T_{X}(\mathrm{pr}_{X}(j_{*}E_{\pm}))\neq\mathrm{pr}_{X}(j_{*}E_{\pm})$. Thus, the result follows. ∎
@@ 3.4. Bayer–Macrì divisors
for any morphism $C\to\mathcal{M}^{Y}_{\sigma_{Y}}(a,b)$ from an integral curve $C$, where $\mathcal{E}_{C}$ is the associated family and $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}$ is the base change of $T_{Y}$ [^5] The functor $T_{Y_{C}}:=T_{Y}\otimes\mathrm{id}_{\mathrm{D}^{\mathrm{b}}(C)}\colon\mathrm{D}^{\mathrm{b}}(Y_{C})\to\mathrm{D}^{\mathrm{b}}(Y_{C})$ is defined by the pull-back of the kernel of the composition $\mathrm{D}^{\mathrm{b}}(Y)\xrightarrow{\mathrm{pr}_{Y}}\mathcal{K}u(Y)\hookrightarrow\mathrm{D}^{\mathrm{b}}(Y)$ to $\mathrm{D}^{\mathrm{b}}(Y_{C}\times_{C}Y_{C})$.. Note that $\mathcal{E}_{C}\in\mathrm{D}^{\mathrm{b}}(X_{C})$ by [8, Lemma 8.3(1)]. From [43, Lemma 2.41], we obtain an isomorphism $p_{Y*}(T_{Y_{C}}(\mathcal{E}_{C}))\cong T_{Y}(p_{Y*}(\mathcal{E}_{C}))$. This implies
@@
as $T_{Y}$ acts trivially on $\mathrm{K}_{\mathrm{num}}(\mathcal{K}u(Y))$ by [25, Lemma 4.10].
@@
then (11) follows from the base change of [25, Equation (35)]. This completes the proof of (6). Finally, the ampleness of $\ell_{\sigma_{Y}}$ follows from the facts that $\ell_{\sigma_{X}}$ is ample (cf. [69, Theorem 3.1]) and the morphism $r$ is finite according to Theorem 3.7. ∎
@@
Similarly, using [25, Equation (40)] instead of [25, Equation (35)], we have the following result for GM fivefolds.
@@ 4.1.2. Atomic objects
Such a vector $\tilde{v}(E)\in\tilde{\mathrm{H}}(X,\mathbb{Q})$ is called an extended Mukai vector of $E$ (cf. [6, Definition 4.16] and [5, Proposition 3.3]). In [54, Section 6], the $1$-dimensional subspace of $\tilde{\mathrm{H}}(X,\mathbb{Q})$ spanned by $\tilde{v}(E)$ is called the LLV line of $E$.
@@
By our assumption, $\mathcal{X}_{b}$ is smooth, hence $\mathcal{X}_{b}\hookrightarrow\mathcal{X}$ is a locally complete intersection. Moreover, $\mathrm{codim}_{\mathcal{Y}}(\mathcal{Y}_{b})=\mathrm{codim}_{\mathcal{X}}(\mathcal{X}_{b})=\dim B$. Then using [43, Corollary 2.27], we obtain
@@
As in [5] and [54, Lemma 6.10], for a sheaf $E\in\operatorname{\mathrm{Coh}}(X)$, we have
@@ 4.2. Atomic sheaves supported on Lagrangian subvarieties
By [54, Theorem 6.24] and [5, Proposition 3.11], each torsion atomic sheaf is supported on points or Lagrangian subvarieties. We make the following definition.
@@
is the projection. Since $\chi_{i_{*}\operatorname{\mathcal{O}}_{L}}$ is of rank one, it follows that $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ has rank at most one for each $t=1,2$. Then the result follows from the non-triviality of $\chi_{i_{*}\operatorname{\mathcal{O}}_{M_{t}}}$ as in the arguments in [54, Lemma 3.9, Remark 3.10] and [5, Section 7.2]. ∎
@@ 4.3. A criterion
To this end, we only need to generalize [5, Proposition 7.2] to our setting. Indeed. the argument there applies analogously without changes. We denote $\mathrm{ker}(\chi^{\mathrm{coh}}_{i_{*}[L]})$ by $K$. Because the image of $i$ is Lagrangian, we see $\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X})\subset K$. Moreover, by the existence of the $\mathfrak{sl}_{2}$-triple $(e_{\omega},h_{\omega},\Lambda_{\omega})\subset\mathfrak{g}(X)_{\mathbb{C}}$ (cf. [72, Section 2]) where $\omega\in\mathrm{H}^{0}(X,\Omega_{X}^{2})$ is the holomorphic symplectic 2-form, the same calculation as in [5, Proposition 7.2] gives $\mathrm{H}^{0}(X,\wedge^{2}\mathcal{T}_{X})\subset K$.
@@
$i_{*}[L]\cup-\colon\mathrm{H}^{2}(X,\mathbb{Q})\xrightarrow{i^{*}}\mathrm{H}^{2}(L,\mathbb{Q})\xrightarrow{i_{*}}\mathrm{H}^{2n+2}(X,\mathbb{Q}).$
@@
The proof of [77, Lemma 1.5] adapts to our situation without changes. It suffices to prove this for cohomology with real coefficients. By the adjunction formula, we have an inclusion $\ker(i^{*})\subset\ker(i_{*}[L]\cup-)$. To show the reverse inclusion, as in the proof of [77, Lemma 1.5], we choose a Kähler class $\lambda\in\mathrm{H}^{2}(X,\mathbb{R})$ such that $i^{*}\lambda$ is also a Kähler class. Now it suffices to prove that the intersection form
@@
is non-degenerate after restricting it to $\mathrm{im}(i^{*})\subset\mathrm{H}^{1,1}_{\mathbb{R}}(L)$. Then the result follows from the Hodge index theorem as in [77, Lemma 1.5] since $i^{*}\lambda$ is a Kähler class. ∎
@@
Let $\omega$ be the holomorphic symplectic $2$-form of $X$. We denote by $\mathcal{N}^{\vee}_{L/X}$ the $-1$ degree cohomology sheaf of the cotangent complex $\mathbb{L}_{L/X}$ of $i$. When $i$ is a closed embedding, $\mathcal{N}^{\vee}_{L/X}$ is exactly the conormal bundle of $L$ in $X$. Since $i$ is a local complete intersection morphism by [71, Tag 0E9K], $\mathcal{N}^{\vee}_{L/X}$ is a bundle (cf. [71, Tag 0FK2] and [71, Tag 068B]). Using the standard exact triangle of cotangent complexes by [71, Tag 0E44], we obtain exact sequences
@@
As $\omega$ is a holomorphic symplectic $2$-form, it induces an isomorphism $\omega\colon\mathcal{T}_{X}\xrightarrow{\cong}\Omega_{X}$. To adapt the argument in [5, Lemma 7.4], it suffices to prove that there exists an isomorphism $\mathcal{N}_{L/X}\cong\Omega_{L}$. According to [33, Remark 4.4], there is a natural map
@@ 5. Atomic sheaves on Bridgeland moduli spaces
In this section, we aim to construct atomic sheaves on $M^{X}_{\sigma_{X}}(a,b)$ for a GM fourfold or sixfold $X$ via two different methods. The main results in this section are Theorem 5.5 and Theorem 5.8. At the end of this section, we prove Theorem 1.2.
@@ 5.2. Fixed loci of involutions are $1$-obstructed
First, we assume that $X$ is non-Hodge-special. As in [67, Section 8.5], up to an equivalence of Kuznetsov components, it is enough to prove the statement when $X$ is a non-Hodge-special special GM fourfold. Let $j\colon Y\hookrightarrow X$ be the branch divisor of $X$, which is an ordinary GM threefold. By [25, Proposition 5.3], for any Serre-invariant stability condition $\sigma_{Y}$ on $\mathcal{K}u(Y)$, the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism
@@
Next, we assume that $X$ is an arbitrary smooth GM fourfold. By [66, Theorem 4.18], we may assume that $X$ is general in the sense that the canonical quadric surface in $X$ is smooth (cf. [66, Remark 2.2]). Therefore, using Theorem 3.2, we can find a smooth connected quasi-projective curve $C$ and a family $\mathcal{X}\to C$ of GM fourfolds with a full triangulated subcategory $\mathcal{K}u(\mathcal{X})\subset\mathrm{D}^{\mathrm{b}}(\mathcal{X})$ such that $\mathcal{K}u(\mathcal{X})_{c}\simeq\mathcal{K}u(\mathcal{X}_{c})$ for each $c\in C$. Additionally,
@@
From [13, Lemma 5.9], there is an auto-equivalence $T_{\mathcal{X}}$ on $\mathcal{K}u(\mathcal{X})$ such that $(T_{\mathcal{X}})|_{c}\simeq T_{\mathcal{X}_{c}}$ for each $c\in C$. By [13, Proposition 5.7] and the proof of [67, Theorem 8.2], such equivalence induces an automorphism $\tau_{\mathcal{X}}\colon M^{\mathcal{X}}_{\underline{\sigma}}(a,b)\to M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$, which is an involution since it is an involution at each fiber over $C$ and $M^{\mathcal{X}}_{\underline{\sigma}}(a,b)$ is smooth. Therefore, the fixed locus $\mathrm{Fix}(\tau_{\mathcal{X}})$ is smooth projective over $C$ and satisfies $(\mathrm{Fix}(\tau_{\mathcal{X}}))|_{c}\cong\mathrm{Fix}(\tau_{\mathcal{X}_{c}})$ for each $c$. By Proposition 5.1, the restriction map
@@
If $Y$ is an ordinary GM fivefold, by [42, Corollary 6.5], we can find an ordinary GM threefold $Y^{\prime}$ with an equivalence $\mathcal{K}u(Y)\simeq\mathcal{K}u(Y^{\prime})$. Thus, we only need to prove the statement when $Y$ is an ordinary GM threefold. As $Y$ is ordinary, it is the branch divisor of a special GM fourfold $X$.
Let $\sigma_{X}\in\operatorname{Stab}^{\circ}(\mathcal{K}u(X))$ be a generic stability condition on $\mathcal{K}u(X)$ with $Z_{\sigma_{X}}(\Lambda_{i})=Z_{\sigma_{Y}}(\lambda_{i})$ for each $i=1,2$. By [25, Proposition 5.3], the functor $\mathrm{pr}_{X}\circ j_{*}$ induces a morphism